The age of father is twice that of the elder son . Ten years hence the age of father will be be three times that of younger son . If the difference between of ages of the two sons is 15 years then the age of father
Answer by ikleyn(52776)

.
The age of father is twice that of the elder son. Ten years hence the age of father will be be three times that of younger son.
If the difference between of ages of the two sons is 15 years then the age of father
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

Let "x" be the present age of the elder son.
Then the father's present age is 2x, while the younger son's present age is (x-15).

In 10 years: The father's age will be      (2x)+10.
             the younger son's age will be (x-15) + 10 = x-5.

From the other side, in ten years the age of the father will be 3(x-5).

Thus you get an equation

2x + 10 = 3(x-5).

Simplify and solve for x:

2x + 10 = 3x - 15  --->  10 + 15 = 3x - 2x  --->  x = 25.

Answer. The elder son's present age is 25 years.
              The younger son's pesent age is 25-15 = 10 years.
              The father's present age is 2*25 = 50 years.


There is a bunch of lessons on age word problems
    - Age problems and their solutions
    - A fresh formulation of a traditional age problem
    - Really intricate age word problem
    - Selected age word problems from the archive
    - Age problem for the day of April, 1
    - OVERVIEW of lessons on age problems
in this site.

Read them and become an expert in solving age problems.

Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Age word problems".